OpenStudy (anonymous):

Assume sqrt(5) is irrational. By using the definition of a rational number, give a proof ny contradiction that for any rational number q, the expression 3q+sqrt(5) is irrational

OpenStudy (chosenmatt):

WHALECOME TO OPENSTUDY

OpenStudy (anonymous):

thanks

OpenStudy (chosenmatt):

lol

OpenStudy (michele_laino):

please, what is q?

OpenStudy (michele_laino):

@addishuns

OpenStudy (anonymous):

i believe it is just any random number, the question does not.

OpenStudy (anonymous):

say what q is*

OpenStudy (anonymous):

I put the whole question, maybe that will help

OpenStudy (michele_laino):

@addishuns Can I suppose q=1?

OpenStudy (anonymous):

i believe so

geerky42 (geerky42):

No, because then you just showed that it is true for q=1. @Michele_Laino You need to show that \(3q+\sqrt5\) is irrational for all \(q\in\mathbb{Q}\) (all rational).

OpenStudy (freckles):

so q=a/b where a,b are integers (not equal to zero) \[\frac{3a}{b}+\sqrt{5}\] Now since you are asking for a proof by contradiction assume \[\frac{3a}{b}+\sqrt{5}=\frac{m}{n}\] where m and n are integers (n not equal to zero) solve for sqrt(5) and see if you can conclude anything from the equation

OpenStudy (anonymous):

\[\sqrt{5} = mb - 3an / bn\]

geerky42 (geerky42):

Right, what do you noticed about \(\dfrac{mb-3an}{bn}\)? *cough* it is rational *cough*

OpenStudy (anonymous):

Of course, so it is a contradiction and proves the original statement.

OpenStudy (perl):

nice job guys